# SAFES: Sequential Privacy and Fairness Enhancing Data Synthesis for Responsible AI

### v1

@@ Abstract
As data-driven and AI-based decision making gains widespread adoption in most disciplines, it is crucial that both data privacy and decision fairness are appropriately addressed. While differential privacy (DP) provides a robust framework for guaranteeing privacy and several widely accepted methods have been proposed for improving fairness, the vast majority of existing literature treats the two concerns independently. For methods that do consider privacy and fairness simultaneously, they often only apply to a specific machine learning task, limiting their generalizability. In response, we introduce SAFES, a Sequential PrivAcy and Fairness Enhancing data Synthesis procedure that sequentially combines DP data synthesis with a fairness-aware data transformation. SAFES allows full control over the privacy-fairness-utility trade-off via tunable privacy and fairness parameters. We illustrate SAFES by combining AIM, a graphical model-based DP data synthesizer, with a popular fairness-aware data pre-processing transformation. Empirical evaluations on the Adult and COMPAS datasets demonstrate that for reasonable privacy loss, SAFES-generated synthetic data achieve significantly improved fairness metrics with relatively low utility loss.
@@ 1 Introduction
Data-driven and AI-based decision making are being adopted in many disciplines and data collected as part of this process often contain sensitive data from individuals. These data are frequently used to make socially impactful decisions including, but not limited to, determining who gets approved for a loan, predicting which applicants should be hired for a job, or forecasting which previously convicted criminals will re-offend. While such applications can and do have legitimate benefits, it is of paramount importance to ensure the use of such sensitive data is responsible and carried out with the highest possible ethical standards.
@@
Since datasets with privacy concerns are likely to also have fairness concerns, and vice versa, it is critical to develop efficient privacy- and fairness-enhancing methods for releasing and analyzing data. In this work, we propose a framework for synthesizing data that simultaneously addresses privacy and fairness concerns by strategically and sequentially combining privacy-preserving data synthesis and fairness-aware pre-processing transformations. We aim to provide a generalized solution that safeguards sensitive personal information, upholds fairness, and keeps the utility of released data close to the original.
@@
There remains a scarcity of research on general-use synthetic data that both satisfies DP guarantees and reduces structural bias, representing a critical area for advancing responsible AI. We address this gap by proposing SAFES – a Sequential PrivAcy and Fairness Enhancing data Synthesis procedure – which combines DP data synthesis with a fairness-aware pre-processing transformation. To our knowledge, this is the first work to attempt such an approach. The output of SAFES is a synthetic dataset that both has theoretical privacy guarantees and has been adjusted in a way to improve structural bias, which in turn improves fairness metrics for downstream classifiers.
@@
SAFES has several benefits. First, it is fully tunable with regards to both the privacy guarantees and fairness constraints. Second, for tight fairness constraints, SAFES exhibits fairness robustness measured by various metrics across a wide range of privacy guarantees per our empirical results, implying that one can adjust the balance between privacy and utility without a significant sacrifice in fairness. Third, though our examples and experiments focus on a commonly used DP data synthesizer and a well-known fairness-aware data transformation methods, SAFES is a general framework and can admit different DP synthesizers of various DP guarantees and different fairness-aware data transformations satisfying various fairness metrics.
2 Related works
2.1 Data synthesis with formal privacy guarantees
@@
Among the concepts developed for privacy protection, differential privacy (DP) (Dwork et al., 2006b, ; Dwork et al., 2006a, ) has emerged as the state-of-the-art framework for guaranteeing privacy when releasing information from sensitive data. To ensure DP, carefully calibrated random noise is added to outputs before release. DP mechanisms generally incur a loss of utility as a trade-off, governed by tunable privacy loss parameters. Many mechanisms have been developed for training ML models (Chaudhuri et al.,, 2011; Abadi et al.,, 2016) that satisfy DP. However, the outputs of these methods are limited to the specific chosen task, and the privacy loss accumulates with each task executed, which limits the number of analyses that can be performed under a fixed privacy budget. This fact motivates the use of DP for synthesizing datasets to be privacy-preserving counterparts to an original, sensitive dataset. Data users can analyze DP-synthesized data as if it were the original data without additional privacy loss due to the immunity to post-processing property of DP. Popular DP data synthesis methods include marginal-based synthesizers for discrete data (Zhang et al.,, 2017; McKenna et al.,, 2019; Eugenio and Liu,, 2021; McKenna et al.,, 2021, 2022). They compute a set of marginals with DP noise and train a model based on the sanitized marginals, from which data are synthesized. Statistical models and deep generative models are also used for DP data synthesis (Zhang et al.,, 2018; Bowen and Liu,, 2020; National Academies of Sciences, Engineering, and Medicine,, 2024).
2.2 Fairness-aware methods
A substantial amount of research and applications on data and algorithmic fairness in the context of AI/ML are for binary classification tasks due to its prevalence of the problem in practice and well-defined fairness metric, among other considerations. We focus on fairness in binary classification in the paper, unless stated otherwise.
@@ 2.2 Fairness-aware methods
Methods for ensuring fairness include pre-processing, in-processing, and post-processing procedures. Pre-processing methods (Hajian and Domingo-Ferrer,, 2013; Calmon et al.,, 2017) transform the dataset to be more fair prior to analysis. In-processing (Kamishima et al.,, 2011; Fish et al.,, 2016) and post-processing (Kamiran et al.,, 2012; Hardt et al.,, 2016) methods, on the other hand, modify the training process and the ultimate model decisions, respectively. While each type of method takes a different approach, the ultimate goal of all of the fairness methods described in this paper is to ensure decisions made by ML classifiers, where these concerns often arise, are done in a fair way. Pre-processing methods include methods that flip the protected attribute and/or outcome label until the proportion of each group receiving each outcome is similar (Hajian and Domingo-Ferrer,, 2013), and methods that use an optimization problem to learn a randomized pre-processing transformation balancing discrimination, distortion, and utility (Calmon et al.,, 2017). The ultimate goal in each case is to remove structural bias from the dataset itself, which in turn improves the fairness of classifiers trained with the pre-processed data. These methods are flexible because they are agnostic to downstream analysis and learning tasks and methods. In-processing methods include incorporating a regularizer when training an ML or AI model to penalize over-reliance on a protected attribute (Kamishima et al.,, 2011) and shifting the classifier decision boundary for unprivileged groups (Fish et al.,, 2016). Post-processing methods include defining a decision rule for a probabilistic classifier that assigns members of the unprivileged group the favorable outcome if the certainty of the classifier’s decision is below a certain threshold (Kamiran et al.,, 2012), and modifying a previously learned classifier to be non-discriminatory by solving a linear program (Hardt et al.,, 2016). Like DP, the methods for achieving fairness also generally incur a loss in utility.
2.3 Interplay between privacy, fairness, and utility
@@
The impacts of privacy on fairness and vice versa have also been studied, including relationship between classification fairness and DP (Dwork et al.,, 2012); whether DP guarantees are inequitably applied across groups, exacerbating unfairness (Ekstrand et al.,, 2018); and whether fair ML techniques can put underprivileged groups at a greater privacy risk (Chang and Shokri,, 2021). It is reported that DP synthetic data often magnify unfairness (Ganev et al.,, 2022) and fairness transformations unevenly distribute privacy risk to underprivileged groups (Chang and Shokri,, 2021). There also exists a three-way trade-off between privacy, fairness, and utility. It is claimed that it is impossible for a single mechanism to achieve both DP and fairness with non-trivial classification accuracy (Cummings et al.,, 2019; Agarwal,, 2021); Classifiers trained on synthesized images via DP generative adversarial networks (GANs) exhibit reduced utility without fairness improvements (Cheng et al.,, 2021).
@@
For specific learning tasks with DP and fairness constraints, methods exist for empirical risk minimization (Ding et al.,, 2020), logistic regression (Xu et al.,, 2019), and stochastic gradient descent (Tran et al.,, 2021), among others. Undersampling a training dataset prior to DP synthesis may also produce better fairness metrics on downstream classification tasks (Bullwinkel et al.,, 2022). Modifications have also been considered to achieve DP for in- and post-processing fairness methods (Jagielski et al.,, 2019). Data synthesis with DP and “justifiable fairness” (Salimi et al.,, 2019) guarantees can be achieved for a marginal-based DP synthesizer (McKenna et al.,, 2021) by ensuring all directed paths between the protected attribute(s) and the response variable in the graphical model representation pass through a non-protected attribute (Pujol et al.,, 2023). however, since the fairness modification is entangled with data synthesis in this approach, it does not apply in the increasingly common scenario where DP synthetic data were already released without fairness considerations. In addition, the approach only achieves justifiable fairness, which is a binary condition (yes or no) and only one of many possible fairness definitions and does not guarantee fairness by other metrics, especially those with a continuous fairness parameter for tuning the trade-off between utility and fairness constraints.
@@ 3 Definitions and notations
We introduce, in this section, the definitions and notations employed in our approach on differential privacy and fairness.
3.1 Differential privacy
Differential privacy (DP) provides a theoretical framework for privacy by bounding the influence of a single individual in a dataset on outputs from the dataset. Let $d(D,D^{\prime})=1$ denote two neighboring datasets $D,D^{\prime}$ differing by one individual.
Definition 1 (($\varepsilon,\delta)$-differential privacy (Dwork et al., 2006a, ; Dwork et al., 2006b, )).
A randomized mechanism $\mathcal{M}$ is ($\varepsilon,\delta)$-differentially private if for all $S\subset\textnormal{Range}(\mathcal{M})$ and for any neighboring datasets $D,D^{\prime}$,
@@ 3.1 Differential privacy
Essentially, DP ensures that a mechanism output cannot differ “too much” based on the inclusion or exclusion of a single individual. The individual differing between the neighboring datasets is arbitrary, meaning that DP guarantees hold simultaneously for all members of the dataset. Smaller $\varepsilon$ corresponds to more privacy. $\delta$ is often interpreted as the probability that $\varepsilon$-DP fails and is usually of $o(1/\mbox{poly}(n))$, where $n$ is the data sample size. There are various extensions (Bun and Steinke,, 2016; Mironov,, 2017; Dong et al.,, 2019) to the original DP definition in Definition 1. We present one of these extensions – zero-Concentrated DP that is used in our experiments.
@@
Let $\mathcal{M}$ be a mechanism satisfying $\rho$-zCDP. For any given $\varepsilon\geq 0$, $\mathcal{M}$ satisfies $(\varepsilon,\delta)$-DP with $\delta=\min_{\alpha>1}\frac{\exp\{(\alpha-1)(\alpha\rho-\varepsilon)\}}{\alpha-1}\left(1-\frac{1}{\alpha}\right)^{\alpha}$.
To release an output from a function on $D$ with DP guarantees, the randomized mechanism $\mathcal{M}$ is calibrated according to the global sensitivity of the function, defined as follows (though originally defined using the $\ell_{1}$-norm (Dwork et al., 2006b, ), we use a more general definition).
Definition 3 ($\ell_{p}$-global sensitivity (GS) (Liu,, 2019)).
Let $f$ be a (potentially vector-valued) function of a dataset $D$. The $\ell_{p}$-GS of $f$ is
@@
The GS of a function is the maximum difference in the function outputs on two neighboring datasets. The larger the GS is, the more noise is needed to achieved a fixed level of privacy guarantee via $\mathcal{M}$. The Gaussian mechanism (Definition 4) and the exponential mechanism (Definition 5) are two commonly used DP mechanisms.
@@
Both $(\varepsilon,\delta)$-DP and $\rho$-zCDP are composable under repeated applications of randomized mechanisms to the same data. Let $\mathcal{D}$ be the space of all possible datasets $D$ and let $\mathcal{M}_{1}:\mathcal{D}\rightarrow\mathcal{R}_{1}$ and $\mathcal{M}_{2}:\mathcal{D}\times\mathcal{R}_{1}\rightarrow\mathcal{R}_{2}$ satisfy $(\varepsilon_{1},\delta_{1})$-DP ($\rho_{1}$-zCDP) and $(\varepsilon_{2},\delta_{2})$-DP ($\rho_{2}$-zCDP), respectively. Then $\mathcal{M}_{2}(D,\mathcal{M}_{1}(D))$ satisfies $(\varepsilon_{1}+\varepsilon_{2},\delta_{1}+\delta_{2})$-DP ($(\rho_{1}+\rho_{2})$-zCDP). For $(\varepsilon,\delta)$-DP, this is known as basic composition (McSherry,, 2009). There exists an advanced composition theorem for DP that gives a tighter privacy loss bound (Dwork et al.,, 2010), but zCDP is still tighter (Bun and Steinke,, 2016). This motivates the practice of specifying the overall privacy loss in terms of $(\varepsilon,\delta)$-DP, applying Theorem 1 to convert it to $\rho$-zCDP, and using DP procedures of $\rho$-zCDP to compose privacy loss.
Another useful property of DP, in addition to privacy loss composition above, is its immunity to post-processing (Dwork et al., 2006b, ; Dwork et al., 2006a, ; Bun and Steinke,, 2016). That is, a post-processing procedure on an output of a DP mechanism, whether satisfying $(\varepsilon,\delta)$-DP or $\rho$-zCDP, does not incur further privacy loss, as long as the procedure does not access the original data. In the case of synthetic data generated from a DP synthesizer, any tasks done on the synthetic data would also satisfy the same DP guarantees as the synthetic data itself.

### v3

@@ Abstract
As data-driven and AI-based decision making gains widespread adoption across disciplines, it is crucial that both data privacy and decision fairness are appropriately addressed. Although differential privacy (DP) provides a robust framework for guaranteeing privacy and methods are available to improve fairness, most prior work treats the two concerns separately. Even though there are existing approaches that consider privacy and fairness simultaneously, they typically focus on a single specific learning task, limiting their generalizability. In response, we introduce SAFES, a Sequential PrivAcy and Fairness Enhancing data Synthesis procedure that sequentially combines DP data synthesis with a fairness-aware data preprocessing step. SAFES allows users flexibility in navigating the privacy-fairness-utility trade-offs. We illustrate SAFES with different DP synthesizers and fairness-aware data preprocessing methods and run extensive experiments on multiple real datasets to examine the privacy-fairness-utility trade-offs of synthetic data generated by SAFES. Empirical evaluations demonstrate that for reasonable privacy loss, SAFES-generated synthetic data can achieve significantly improved fairness metrics with relatively low utility loss.
@@ 1 Introduction
1.1 Background
Data-driven and AI-based decision making is increasingly adopted across disciplines. These systems are frequently used to make socially consequential decisions including, but not limited to, loan approvals, hiring, and criminal justice decisions. Yet the data collected in this process often contain sensitive personal information. While such AI applications can and do have legitimate benefits, it is of paramount importance to ensure sensitive data are used responsibly and in accordance with the highest possible ethical standards.
@@ 1.1 Background
Since datasets with privacy concerns are likely to also have fairness concerns and vice versa, it is critical to develop efficient privacy- and fairness-enhancing methods for releasing and analyzing data. In this work, we propose a framework for synthesizing data that simultaneously addresses privacy and fairness concerns by strategically and sequentially combining privacy-preserving data synthesis and fairness-aware data preprocessing. We aim to provide a generalized solution that safeguards sensitive personal information, upholds fairness, and keeps the utility of released data close to the original simultaneously.
@@
There remains a scarcity of research on general-use synthetic data that both satisfies DP guarantees and reduces structural bias, representing a critical area for advancing responsible AI. We address this gap by proposing SAFES – a Sequential PrivAcy and Fairness Enhancing data Synthesis procedure – which combines DP data synthesis with a fairness-aware data preprocessing. The output of SAFES is a synthetic dataset with formal privacy guarantees and improved fairness for downstream classifiers, largely preserved utility for general and ML applications.
@@
While applied in focus, this work is—to the best of our knowledge—the first conceptual framework to integrate differential privacy and fairness sequentially for synthetic data generation, and the first to empirically assess the framework’s utility and feasibility. We believe SAFES has several benefits. First, it is flexible with respect to both the level of privacy protection and the degree of fairness enforced (with certain data preprocessing methods for fairness). Second, SAFES exhibits fairness robustness measured by various metrics across a wide range of privacy guarantees per our empirical results, implying that one can adjust the balance between privacy and utility without a significant sacrifice in fairness. Third, SAFES is a general framework that accommodates diverse DP synthesizers (with various types of DP privacy guarantees) and fairness-aware preprocessing methods aligned with specified fairness metrics.
1.2 Related works
Data synthesis with formal privacy guarantees
@@ 1.2 Related works
Among the concepts developed for privacy protection, differential privacy (DP) (Dwork et al., 2006b, ; Dwork et al., 2006a, ) has emerged as the state-of-the-art framework for guaranteeing privacy when releasing information from sensitive data. To ensure DP, calibrated random noise at given privacy loss parameters is added to the outputs via DP mechanisms before release. Many mechanisms have been developed for executing queries (Hardt et al.,, 2012; Kotsogiannis et al.,, 2019) and training ML models (Chaudhuri et al.,, 2011; Abadi et al., 2016a, ) that satisfy DP. However, the outputs of these methods are limited to the specific chosen task, and the privacy loss accumulates with each task executed, which limits the number of analyses that can be performed under a fixed privacy budget. This fact motivates the use of DP for synthesizing datasets to be privacy-preserving counterparts to an original, sensitive dataset. Data users can analyze DP-synthesized data as if it were the original data without additional privacy loss due to the immunity to post-processing property of DP. Popular DP data synthesis methods include marginal-based synthesizers for discrete data (Zhang et al.,, 2017; McKenna et al.,, 2019; Eugenio and Liu,, 2021; McKenna et al.,, 2021; McKenna et al., 2022a, ). They compute a set of marginals with DP noise and train a model based on the sanitized marginals, from which data are synthesized. Statistical models and deep generative models are also used for DP data synthesis (Zhang et al.,, 2018; Bowen and Liu,, 2020; National Academies of Sciences, Engineering, and Medicine,, 2024).
Fairness-aware methods
@@
We focus on fairness for binary classification tasks in this paper, unless otherwise stated. This is because most work – in the literature and in practice – on fairness in AI/ML has focused on binary classification, since fairness challenges often appear in such problems. As a result, this area is the most studied and practically relevant. Methods for ensuring fairness include preprocessing, in-processing, and post-processing procedures. processing methods (Kamiran and Calders,, 2012; Hajian and Domingo-Ferrer,, 2013; Calmon et al.,, 2017) transform the dataset to be more fair prior to analysis. In-processing (Kamishima et al.,, 2011; Fish et al.,, 2016) and post-processing (Kamiran et al.,, 2012; Hardt et al.,, 2016) methods, on the other hand, modify the training process and the ultimate model decisions, respectively. While each type of method takes a different approach, the ultimate goal is to ensure decisions made by ML classifiers are fair. Preprocessing methods aims to remove structural bias from the dataset itself, which in turn improves the fairness of classifiers trained with the preprocessed data. Specifically, those methods may assign weights to different dataset records (Kamiran and Calders,, 2012), flip the protected attribute and/or outcome label until the proportion of each group receiving each outcome is similar (Hajian and Domingo-Ferrer,, 2013), and learn a randomized pre-processsor balancing discrimination, distortion, and utility (Calmon et al.,, 2017). The methods are flexible because they are agnostic to downstream analysis and learning tasks. In-processing methods may incorporate regularizers when training a model to penalize over-reliance on a protected attribute (Kamishima et al.,, 2011) or shift the classifier decision boundary for unprivileged groups (Fish et al.,, 2016). Post-processing approaches include adjusting decision rules for probabilistic classifiers to favor unprivileged groups when prediction confidence is low (Kamiran et al.,, 2012), and reformulating trained classifiers to remove discrimination through linear programming (Hardt et al.,, 2016).
Interplay between privacy, fairness, and utility
@@
The mutual influence between privacy and fairness has been widely studied. Research has explored the connection between classification fairness and DP, the potential inequitable application of DP guarantees across demographic groups (Ekstrand et al.,, 2018), and the possibility that fair ML methods increase privacy risks for underprivileged groups (Chang and Shokri,, 2021). Studies report that DP-generated synthetic data can exacerbate unfairness (Ganev et al.,, 2022), and that fairness adjustments may unevenly allocate privacy risks (Chang and Shokri,, 2021). Incorporating DP and fairness in ML/AI models generally entail a reduction in utility that is well-documented in the literature. There also exists a three-way trade-off between privacy, fairness, and utility. (Cummings et al.,, 2019; Agarwal,, 2021) claim that it is impossible for a single mechanism to achieve both DP and fairness with non-trivial classification accuracy. Classifiers trained on synthesized images via DP generative adversarial networks (GANs) exhibit reduced utility without fairness improvements (Cheng et al.,, 2021).
@@
For specific ML tasks with both DP and fairness constraints, methods exist for empirical risk minimization (Ding et al.,, 2020), logistic regression (Xu et al.,, 2019), and stochastic gradient descent (SGD) (Tran et al.,, 2021), among others. In- and post-processing fairness methods may also be modified to achieve DP (Jagielski et al.,, 2019). However, the output of these methods is a specific classifier, unlike our work that produces a full privacy-preserving and fairness-aware individual-level synthetic dataset agnostic to potential downstream uses. (Zhao et al.,, 2025) finds that common post-processing in DP outputs such as non-negativity or preservation of total population counts incur unfairness and varying the privacy budget across different groups mitigates these fairness issues. Under-sampling a training dataset prior to DP synthesis may also produce better fairness metrics on downstream classification tasks (Bullwinkel et al.,, 2022). Data synthesis with DP and “justifiable fairness” (Salimi et al.,, 2019) can be achieved for a marginal-based DP synthesizer (McKenna et al.,, 2021) by ensuring all directed paths between the protected attribute(s) and the response variable in the graphical model representation pass through a non-protected attribute (Pujol et al.,, 2023). However, since the fairness modification is entangled with data synthesis in this approach, it does not apply to scenarios where DP synthetic data were already released without fairness considerations. In addition, it only achieves justifiable fairness – a binary condition (yes or no) and only one of many possible fairness definitions – and does not guarantee fairness by other metrics, especially those with a continuous fairness parameter for tuning the trade-off between utility and fairness constraints.
@@ 2 Definitions and notations
We introduce, in this section, the definitions and notations employed in SAFES and the experiments.
2.1 Differential privacy (DP)
DP provides a theoretical framework for privacy by bounding the influence of a single individual in a dataset on outputs from the dataset. Let $d(D,D^{\prime})=1$ denote two neighboring datasets $D,D^{\prime}$, which differ by one individual.
Definition 1 $($$\varepsilon,\delta)$-differential privacy Dwork et al., 2006a ; Dwork et al., 2006b ).
A randomized mechanism $\mathcal{M}$ is said to be ($\varepsilon,\delta)$-differentially private if for all $S\subset\textnormal{Range}(\mathcal{M})$ and for any neighboring datasets $D,D^{\prime}$,
@@ 2.1 Differential privacy (DP)
Essentially, DP ensures that a mechanism output cannot differ too much with vs.  without any single individual. Smaller $\varepsilon$ corresponds to more privacy. $\delta$ is often interpreted as the probability that $\varepsilon$-DP fails and is usually of $o(1/\mbox{poly}(n))$, where $n$ is the data sample size. There are various extensions (Bun and Steinke,, 2016; Mironov,, 2017; Dong et al.,, 2019) to the original DP definition in Definition 1. We present one of these extensions – zero-concentrated DP – that is used in our experiments.
@@
Let $\mathcal{M}$ be a mechanism satisfying $\rho$-zCDP. For any given $\varepsilon\geq 0$, $\mathcal{M}$ satisfies $(\varepsilon,\delta)$-DP with
$\delta=\min_{\alpha>1}\frac{\exp\{(\alpha-1)(\alpha\rho-\varepsilon)\}}{\alpha-1}\left(1-\frac{1}{\alpha}\right)^{\alpha}.$
To release an output from a function on $D$ with DP guarantees, mechanism $\mathcal{M}$ is calibrated according to the global sensitivity (GS) of the function, defined as follows (though originally defined using the $\ell_{1}$-norm (Dwork et al., 2006b, ), we use a more general definition $\ell_{p}$-GS (Liu,, 2019). Let $f$ be a (potentially vector-valued) function of a dataset $D$. The $\ell_{p}$-GS of $f$ is
@@
The GS of a function is the maximum difference in the function outputs on two neighboring datasets. The larger the GS is, the more noise is needed to achieve a fixed level of privacy guarantee via $\mathcal{M}$. The Gaussian mechanism (Definition 4) and the exponential mechanism (Definition 5) are two commonly used DP mechanisms.
@@
Both $(\varepsilon,\delta)$-DP and $\rho$-zCDP are composable under repeated applications of randomized mechanisms to the same data. Let $\mathcal{D}$ be the space of all possible datasets $D$ and let $\mathcal{M}_{1}:\mathcal{D}\rightarrow\mathcal{R}_{1}$ and $\mathcal{M}_{2}:\mathcal{D}\times\mathcal{R}_{1}\rightarrow\mathcal{R}_{2}$ satisfy $(\varepsilon_{1},\delta_{1})$-DP ($\rho_{1}$-zCDP) and $(\varepsilon_{2},\delta_{2})$-DP ($\rho_{2}$-zCDP), respectively. Then $\mathcal{M}_{2}(D,\mathcal{M}_{1}(D))$ satisfies $(\varepsilon_{1}+\varepsilon_{2},\delta_{1}+\delta_{2})$-DP ($(\rho_{1}+\rho_{2})$-zCDP). For $(\varepsilon,\delta)$-DP, this is known as basic composition (McSherry,, 2009). While advanced composition gives a tighter privacy loss bound for $(\varepsilon,\delta)$-DP (Dwork et al.,, 2010), zCDP still provides tighter privacy loss bounds (Bun and Steinke,, 2016; Mironov,, 2017). This motivates the practice of specifying the overall privacy loss in terms of $(\varepsilon,\delta)$-DP, applying the conversion theorems to convert it to $\rho$-zCDP, then using zCDP procedures to compose privacy loss.
Another useful property of DP, in addition to the privacy loss composition described above, is its immunity to post-processing (Dwork et al., 2006b, ; Dwork et al., 2006a, ; Bun and Steinke,, 2016). That is, a post-processing procedure on an output of a DP mechanism, whether satisfying $(\varepsilon,\delta)$-DP or $\rho$-zCDP, does not incur further privacy loss, as long as the procedure does not access the original data. In the case of synthetic data generated from a DP synthesizer, any tasks performed on the synthetic data would satisfy the same DP guarantees as the synthetic data itself.